OpenStudy (anonymous):

If f(x) = 2x +4, g(x) = x^2-1 and h(x) = square root x+4, what is (f o g o h)(4)?

OpenStudy (anonymous):

goh = (sqrt(x+4))^2 -1 = x+4-1=x+3 , f o g o h = 2(x+3) +4 = 2x+10 ===> f o g o h (4) = 2*4+10 = 18